# Driving on Registers

### v1

@@ 4.2.1 Perception
Various previous methods [feng2025rap, guo2025ipad] use pretrained backbones. We first study the impact of such an initialization on the planning results in 4(a). First, a good initialization of the perception ViT is crucial: using a pretrained backbone improves scores by a large margin (more than 15 PDMS points). Second, using large-scale pretrained DINOv2 [oquab2024dinov2] improves further over pretraining on ImageNet21k [kolesnikov2020big]. In all subsequent experiments, we use a pretrained DINOv2 backbone.
@@
As argued in section 3, using few perceptual tokens is of interest as it makes the trajectory prediction and scoring lightweight and faster. In 4(b), first column, we study three compression approaches: using a pooling operation of the output feature map as in [li2025generalized] (c), using a transformer decoder with 16 queries per camera (e), and our proposed approach using 16 additional registers in the model (h). For comparison purpose, we also provide models using the full feature maps, i.e., 16k scene tokens in (b). We observe that our register-based approach outperforms the pooling operation while introducing a low-overhead (0.6M parameters) at training. It also nearly reaches the performances of the no-compression model, despite using 250 times fewer tokens for downstream processing. Additionally, we observe that using a transformer decoder (with roughly the same number of parameters) does not reach the same performance.
@@
Next, we study the finetuning strategy in 4(b), second column, with three different settings: full finetuning (d, f), frozen backbone (a, d, g) and LoRA finetuning [hu2022lora] (b, c, e, h). For frozen backbone with registers, only the registers are learnable. First, we observe that LoRA finetuning improves the results by a large margin, reaching similar conclusion as in [barin2024robust] over frozen backbone. It is also the case when comparing LoRA (h) and Full finetuning (f), but to our understanding, it should be possible to close this gap with careful learning rate scheduling specific to the backbone. LoRA being more robust to these meta parameters, we use LoRA finetuning as default.
@@ 5 Conclusion
We present DrivoR, a novel E2E driving method using register-based compression and disentangled scoring representations. DrivoR highlights that full-transformer architectures without complex intermediate states nor large trajectory dictionaries can achieve state-of-the-art results. Future works may explore compression incorporating historical frames or additional sensors, or exploring data and model scaling effects.
@@ Acknowledgments
We thank Loick Chambon for constant support throughout the project and Lan Feng for helpful discussions. This work was granted access to the HPC resources of IDRIS under the allocations AD011016241 and AD011016239R1 made by GENCI. We acknowledge EuroHPC Joint Undertaking for awarding the project ID EHPC-REG-2024R02-210 access to Karolina, Czech Republic. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. This work was supported by the European Union’s Horizon Europe research and innovation programme under grant agreement No 101214398 (ELLIOT).
[…]
@@ Appendix C Inference weights of the sub-scores
The subscore prediction design in DrivoR enables flexible weighting strategy. In Table 10, we provide the sub-score weights for both PDMS (NAVSIM-v1) and EPDMS (NAVSIM-v2). For simplicity, the model trained for NAVSIM-v1 keeps the same weights as in the standard PDMS metrics [dauner2024navsim]. For NAVSIM-v2, we adjust the weights with the exception of the Comfort. The adjustment is validated with NAVSIM-v2 warmup-two-stage, see Figure 6 in the original paper.
@@
Method Encoder Parameters $\downarrow$ GFLOPs $\downarrow$ Memory $\downarrow$ (ms) $\downarrow$ EPDMS $\uparrow$
RAP-Dino [feng2025rap] ViT-H  [^8] M 4760 4.2GB 690ms 39.6
@@
+GTRS-DP [li2025generalized] +V2-99 +116M +731 +0.75GB +389ms -
GTRS-D [li2025generalized] V2-99 81M 404 1.0GB 96ms 45.0
GTRS-A [li2025generalized] V2-99 171M 439 1.42GB 243ms 45.4
GTRS-D [li2025generalized] ViT-L 321M 1730 1.6GB 400ms 47.0
ZTRS [li2025ztrs] V2-99 81M 840 1.0GB 193ms 48.1
DrivoR ViT-S 34M 341 0.5GB 110ms 48.3
@@
Table 11: Efficiency. We compare the number of parameters, GFLOPs, peak memory consumption, and throughput w.r.t. the NAVSIM-v2 EPDMS performance. GTRS-A refers to GTRS-Aug and GTRS-D for GTRS-Dense.
@@ Appendix F Single- vs multi- token per trajectory
Previous works such as [guo2025ipad] have used a single token per trajectory pose, decoding each token into an $(x,y,\theta)$ tuple. We ablate the choice of decoding a single token to a full trajectory vs decoding a set of tokens. Table 12 shows that mapping a single token to a trajectory leads to a large jump in performance. Figure 7 visualizes the same scene, with trajectories decoded either from multi- or single-token. The trajectories mapped from a single token are much smoother and contain less noise, showing that representing trajectories with a single token simplifies learning.
@@
PDM‑Closed [dauner2023pdm] PMLR’23 - 94.6 99.8 89.9 86.9 99.9 89.1
Human driver [dauner2024navsim] NeurIPS’24 - 100 100 100 99.9 87.5 94.8
@@
Centaur [sima2025centaur] arXiv’25 C 99.5 98.9 98.0 100 85.9 92.6
@@
AutoVLA [zhou2025autovla] NeurIPS’25 C 99.1 97.1 97.1 100.0 87.6 92.1
DriveVLA-W0 [li2025drivevla] arXiv’25 C 99.3 97.4 97.0 99.9 88.3 93.0
TransDiffuser [jiang2025transdiffuser] arXiv’25 C+L 99.4 96.5 97.8 99.4 94.1 94.9
@@
TransFuser [chitta2022transfuser] TPAMI’22 C+L 97.7 92.8 92.8 100 79.2 84.0
DistillDrive [yu2025distilldrive] arXiv’25 C+L 98.1 94.6 93.6 100 81.0 86.2
TrajHF (EM) [li2025learningpersonalizeddrivingstyles] arXiv’25 C+L 96.6 96.6 92.1 100 84.5 87.6
DiffusionDrive [liao2025diffusiondrive] CVPR’25 C+L 98.2 96.2 94.7 100 82.2 88.1
WOTE [li2025end] ICCV’25 C+L 98.5 96.8 94.9 99.9 81.9 88.3
Hydra-MDP (V2-99) [li2024hydra] arXiv’24 C+L 98.0 97.8 93.9 100 86.5 90.3
GoalFlow (V2-99) [xing2025goalflow] CVPR’25 C+L 98.4 98.3 94.6 100 85.0 90.3
ResAD (V2-99) [zheng2025resad] arXiv’25 C+L 98.9 97.8 94.9 100 87.0 90.6
SeerDrive (V2-99) [zhang2025future] NeurIPS’25 C+L 98.8 98.6 95.8 100 84.2 90.7
@@
Ego‑stat. MLP [dauner2024navsim] NeurIPS’24 C 93.0 77.3 83.6 100 62.8 65.6
UniVLA [wang2025unified] arXiv’25 C 96.9 91.1 91.7 96.7 76.8 81.7
DrivingGPT [chen2024drivinggpt] ICCV’24 C 98.9 90.7 94.9 95.6 79.7 82.4
UniAD [hu2023uniad] CVPR’23 C 97.8 91.9 92.9 100 78.8 83.4
LTF [chitta2022transfuser] TPAMI’22 C 97.4 92.8 92.4 100 79.0 83.8
PARA‑Drive [paradrive] CVPR’24 C 97.9 92.4 93.0 99.8 79.3 84.0
DriveX-S [shi2025drivex] ICCV’25 C 97.5 94.0 93.0 100 79.7 84.5
World4Drive [zheng2025world4drive] ICCV’25 C 97.4 94.3 92.8 100 79.9 85.1
DRAMA [yuan2024drama] ISRR’24 C 98.0 93.1 94.8 100 80.1 85.5
VAD-v2 [chen2024vadv2] arXiv’24 C 98.1 94.8 94.3 100 80.6 86.2
PRIX [wozniak2025prix] arXiv’25 C 98.1 96.3 94.1 100 82.3 87.8
DiffusionDrive [liao2025diffusiondrive] CVPR’25 C 98.2 96.2 94.7 100 82.2 88.1
DIVER [song2025breaking] arXiv’25 C 98.5 96.5 94.9 100 82.6 88.3
AutoVLA [zhou2025autovla] NeurIPS’25 C 98.4 95.6 98.0 99.9 81.9 89.1
DriveVLA-W0 [li2025drivevla] arXiv’25 C 98.7 99.1 95.3 99.3 83.3 90.2
ReCogDrive [li2025recogdrive] arXiv’25 C 97.9 97.3 94.9 100 87.3 90.8
Hydra-MDP++ [li2025hydramdppp] arXiv’25 C 98.6 98.6 95.1 100 85.7 91.0
iPad [guo2025ipad] arXiv’25 C 98.6 98.3 94.9 100 88.0 91.7
DriveSuprim [yao2025drivesuprim] arXiv’25 C 98.6 98.6 95.5 100 91.3 93.5
DrivoR (ours) C 99.0 98.9 96.7 100 90.0 93.7
@@ Appendix G Expanded NAVSIM-v1 results
We provide in Table 13 a full comparison to state-of-the-art methods, including methods with LiDAR inputs and with post-processing such test-time training, Best-of-N scoring and ensembling. We note that the Best-of-N practice aims to score all predicted trajectories with ground truth and select the best trajectory based on the scores from the evaluation, which is not aligned with the evaluation protocol defined in NAVSIM [dauner2024navsim, Cao2025navsim2].
@@ Appendix H NAVSIM-v2 results before benchmark fix
From the table, we see that DrivoR ranks among the best-performing methods, with a much lighter model design (see more detail in Appendix E) and without extra training data. Also, DrivoR ranks first among state-of-the-art methods after the bug fix. This indicates that the errors made by DrivoR are mostly due to human errors, which are not penalized after the fix according to the NAVSIM-v2 benchmark [Cao2025navsim2].
@@
PDM-Closed [dauner2023pdm] 94.4 98.8 100 99.5 100 93.5 99.3 97.7 36.0 88.1 90.6 96.3 98.5 100 83.1 73.7 91.5 25.4 51.3
Const. Vel. [dauner2024navsim] 88.8 42.8 70.6 99.3 77.5 87.3 78.6 97.1 60.4 83.2 59.1 76.5 98.0 71.3 81.1 47.9 97.1 61.9 10.9
Ego Hist. MLP [dauner2024navsim] 93.2 55.7 86.6 99.3 81.2 92.2 83.5 97.5 77.7 77.2 51.9 74.4 98.2 77.1 75.0 40.8 97.8 79.8 12.7
LTF [chitta2022transfuser] 96.2 79.6 99.1 99.6 84.1 95.1 94.2 97.6 79.1 77.8 70.2 84.3 98.1 85.1 75.7 45.4 95.8 76.0 23.1
RAP-DINO (ViT-H) [feng2025rap] 97.1 94.4 98.8 99.8 83.9 96.9 94.7 96.4 66.2 83.2 83.9 87.4 98.0 86.9 80.4 52.3 95.2 52.4 36.9
GTRS-D (V2-99) [li2025generalized] 98.7 91.4 95.8 89.2 99.4 94.4 99.3 98.8 72.8 69.5 98.7 90.1 95.1 54.6 96.9 94.1 40.4 49.7 41.7
GTRS-A (V2-99) [li2025generalized] 98.9 87.9 95.1 88.8 99.2 89.6 99.6 98.8 76.1 80.3 99.1 86.0 94.7 53.5 97.6 97.1 54.2 56.1 42.1
@@
GTRS-D EVA-ViT-L [li2025generalized] 97.6 95.8 99.8 99.0 77.2 97.8 95.3 97.3 46.7 91.9 91.3 92.7 99.0 72.7 90.4 53.8 94.1 41.6 43.4
GTRS-A (ViT-L) [li2025generalized] 98.7 98.0 99.1 99.8 75.9 98.7 94.7 97.6 49.8 89.5 89.6 92.9 98.5 78.9 86.4 55.3 96.5 52.7 44.7
DriveSuprim (EVA-ViT-L) [yao2025drivesuprim] 98.7 98.0 99.1 99.8 75.9 98.7 94.7 97.6 49.8 89.5 89.6 92.9 98.5 78.9 86.4 55.3 96.5 52.7 44.7
GTRS-D (ViT-L) [li2025generalized] 98.9 98.2 99.8 99.6 73.9 98.9 95.3 97.3 40.0 91.5 90.8 94.7 98.5 70.8 90.1 55.4 97.2 54.2 45.3
DrivoR (ViT-S) (ours) 98.8 95.1 98.9 100 72.6 98.7 94.0 97.6 73.3 90.2 88.4 91.9 98.6 69.8 88.0 50.1 98.5 76.2 45.3
ZTRS (V2-99) [li2025ztrs] 98.9 97.6 100.0 100.0 66.7 98.9 96.2 96.7 44.0 91.1 90.4 95.8 99.0 63.6 89.8 60.4 97.6 66.1 45.5
@@
RAP-DINO (ViT-H) [feng2025rap] 97.1 94.4 98.8 99.8 83.9 96.9 94.7 96.4 66.2 83.2 83.9 87.4 98.0 86.9 80.4 52.3 95.2 52.4 39.6
GTRS-D (V2-99) [li2025generalized] 98.9 96.2 99.4 99.3 72.9 98.9 95.1 96.9 39.1 91.2 89.4 94.4 98.8 69.5 90.0 54.3 94.0 48.7 45.0
GTRS-A (V2-99) [li2025generalized] 98.9 95.1 99.1 99.6 76.2 99.1 94.9 97.6 54.2 88.1 88.8 89.3 98.9 98.9 85.9 53.7 96.8 56.9 45.4
@@
ZTRS (V2-99) [li2025ztrs] 98.9 97.6 100 100 66.7 98.9 96.2 96.7 44.0 91.1 90.4 95.8 99.0 63.6 89.8 60.4 97.6 66.1 48.1
DrivoR (ViT-S) (ours) 98.8 95.1 98.9 100 72.6 98.7 94.0 97.6 73.3 90.2 88.4 91.9 98.6 70.0 88.0 50.1 98.5 76.2 48.3

### v2

@@ 4.2.1 Perception
Various previous methods [15, 16] use pretrained backbones. We first study the impact of such an initialization on the planning results in 4(a). First, a good initialization of the perception ViT is crucial: using a pretrained backbone improves scores by a large margin (more than 15 PDMS points). Second, using large-scale pretrained DINOv2 [39] improves further over pretraining on ImageNet21k [26]. In all subsequent experiments, we use a pretrained DINOv2 backbone.
@@
As argued in section 3, using few perceptual tokens is of interest as it makes the trajectory prediction and scoring lightweight and faster. In 4(b), first column, we study three compression approaches: using a pooling operation of the output feature map as in [35] (c), using a transformer decoder with 16 queries per camera (e), and our proposed approach using 16 additional registers in the model (h). For comparison purpose, we also provide models using the full feature maps, i.e., 16k scene tokens in (b). We observe that our register-based approach outperforms the pooling operation while introducing a low-overhead (0.6M parameters) at training. It also nearly reaches the performances of the no-compression model, despite using 250 times fewer tokens for downstream processing. Additionally, we observe that using a transformer decoder (with roughly the same number of parameters) does not reach the same performance.
@@
Next, we study the finetuning strategy in 4(b), second column, with three different settings: full finetuning (d, f), frozen backbone (a, d, g) and LoRA finetuning [18] (b, c, e, h). For frozen backbone with registers, only the registers are learnable. First, we observe that LoRA finetuning improves the results by a large margin, reaching similar conclusion as in [1] over frozen backbone. It is also the case when comparing LoRA (h) and Full finetuning (f), but to our understanding, it should be possible to close this gap with careful learning rate scheduling specific to the backbone. LoRA being more robust to these meta parameters, we use LoRA finetuning as default.
@@ 5 Conclusion
We present DrivoR, a novel E2E driving method using register-based compression and disentangled scoring representations. DrivoR highlights that full-transformer architectures without complex intermediate states nor large trajectory dictionaries can achieve state-of-the-art results. Future works may explore compression incorporating historical frames, additional sensors or map information.
@@ Acknowledgments
We thank Loick Chambon for constant support throughout the project and Lan Feng for helpful discussions. This work was granted access to the HPC resources of IDRIS under the allocations AD011016241, AD011016239R1 and AD011012883R4 made by GENCI. We acknowledge EuroHPC Joint Undertaking for awarding the project ID EHPC-REG-2024R02-210 access to Karolina, Czech Republic. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. This work was supported by the European Union’s Horizon Europe research and innovation programme under grant agreement No 101214398 (ELLIOT).
@@ Appendix C Inference weights of the sub-scores
The subscore prediction design in DrivoR enables flexible weighting strategy. In Table 10, we provide the sub-score weights for both PDMS (NAVSIM-v1) and EPDMS (NAVSIM-v2). For simplicity, the model trained for NAVSIM-v1 keeps the same weights as in the standard PDMS metrics [12]. For NAVSIM-v2, we adjust the weights with the exception of the Comfort. The adjustment is validated with NAVSIM-v2 warmup-two-stage, see Figure 6 in the original paper.
@@
Method Img. size Cams Encoder Parameters $\downarrow$ GFLOPs $\downarrow$ Memory $\downarrow$ (ms) $\downarrow$ EPDMS $\uparrow$
RAP-Dino (448, 768) 4 ViT-H  [^8] M 4760 4.2GB 690ms 39.6
[…]
@@
+GTRS-DP (512, 2048) 4 +V2-99 +116M +1249 +1.15GB +389ms -
@@
GTRS-D (512, 2048) 1 V2-99 81M 404 1.0GB 96ms 45.0
[…]
GTRS-A (512, 2048) 1 V2-99 171M 439 1.42GB 243ms 45.4
[…]
GTRS-D (512, 2048) 1 ViT-L 321M 1730 1.6GB 400ms 47.0
[…]
GTRS-D (512, 2048) 1 ViT-S 32M 234 0.38GB 303ms 45.9
[…]
ZTRS (512, 2048) 2 V2-99 81M 840 1.0GB 193ms 48.1
[…]
DrivoR (672, 1148) 4 ViT-S 41M 351 0.5GB 110ms 48.3
[…]
@@ Appendix F Single- vs multi- token per trajectory
Previous works such as [16] have used a single token per trajectory pose, decoding each token into an $(x,y,\theta)$ tuple. We ablate the choice of decoding a single token to a full trajectory vs decoding a set of tokens. Table 12 shows that mapping a single token to a trajectory leads to a large jump in performance. Figure 7 visualizes the same scene, with trajectories decoded either from multi- or single-token. The trajectories mapped from a single token are much smoother and contain less noise, showing that representing trajectories with a single token simplifies learning.
@@
PDM‑Closed [11] PMLR’23 - 94.6 99.8 89.9 86.9 99.9 89.1
Human driver [12] NeurIPS’24 - 100 100 100 99.9 87.5 94.8
@@
Centaur [42] arXiv’25 C 99.5 98.9 98.0 100 85.9 92.6
@@
AutoVLA [61] NeurIPS’25 C 99.1 97.1 97.1 100.0 87.6 92.1
DriveVLA-W0 [30] arXiv’25 C 99.3 97.4 97.0 99.9 88.3 93.0
TransDiffuser [25] arXiv’25 C+L 99.4 96.5 97.8 99.4 94.1 94.9
@@
TransFuser [8] TPAMI’22 C+L 97.7 92.8 92.8 100 79.2 84.0
DistillDrive [55] arXiv’25 C+L 98.1 94.6 93.6 100 81.0 86.2
TrajHF (EM) [28] arXiv’25 C+L 96.6 96.6 92.1 100 84.5 87.6
DiffusionDrive [36] CVPR’25 C+L 98.2 96.2 94.7 100 82.2 88.1
WOTE [31] ICCV’25 C+L 98.5 96.8 94.9 99.9 81.9 88.3
Hydra-MDP (V2-99) [33] arXiv’24 C+L 98.0 97.8 93.9 100 86.5 90.3
GoalFlow (V2-99) [52] CVPR’25 C+L 98.4 98.3 94.6 100 85.0 90.3
ResAD (V2-99) [59] arXiv’25 C+L 98.9 97.8 94.9 100 87.0 90.6
SeerDrive (V2-99) [57] NeurIPS’25 C+L 98.8 98.6 95.8 100 84.2 90.7
@@
Ego‑stat. MLP [12] NeurIPS’24 C 93.0 77.3 83.6 100 62.8 65.6
UniVLA [48] arXiv’25 C 96.9 91.1 91.7 96.7 76.8 81.7
DrivingGPT [7] ICCV’24 C 98.9 90.7 94.9 95.6 79.7 82.4
UniAD [19] CVPR’23 C 97.8 91.9 92.9 100 78.8 83.4
LTF [8] TPAMI’22 C 97.4 92.8 92.4 100 79.0 83.8
PARA‑Drive [49] CVPR’24 C 97.9 92.4 93.0 99.8 79.3 84.0
DriveX-S [41] ICCV’25 C 97.5 94.0 93.0 100 79.7 84.5
World4Drive [58] ICCV’25 C 97.4 94.3 92.8 100 79.9 85.1
DRAMA [56] ISRR’24 C 98.0 93.1 94.8 100 80.1 85.5
VAD-v2 [6] arXiv’24 C 98.1 94.8 94.3 100 80.6 86.2
PRIX [50] arXiv’25 C 98.1 96.3 94.1 100 82.3 87.8
DiffusionDrive [36] CVPR’25 C 98.2 96.2 94.7 100 82.2 88.1
DIVER [43] arXiv’25 C 98.5 96.5 94.9 100 82.6 88.3
AutoVLA [61] NeurIPS’25 C 98.4 95.6 98.0 99.9 81.9 89.1
DriveVLA-W0 [30] arXiv’25 C 98.7 99.1 95.3 99.3 83.3 90.2
ReCogDrive [32] arXiv’25 C 97.9 97.3 94.9 100 87.3 90.8
Hydra-MDP++ [29] arXiv’25 C 98.6 98.6 95.1 100 85.7 91.0
iPad [16] arXiv’25 C 98.6 98.3 94.9 100 88.0 91.7
DriveSuprim [53] arXiv’25 C 98.6 98.6 95.5 100 91.3 93.5
DrivoR (train) C 98.9 98.3 96.2 100 89.1 93.1
DrivoR (trainval) C 99.0 98.9 96.7 100 90.0 93.7
DrivoR (+65k SimScale data) C 99.1 99.0 96.9 100 90.3 94.0
DrivoR (+134k SimScale data) C 99.1 99.2 96.9 100 91.6 94.6
@@ Appendix G Expanded NAVSIM-v1 results
We provide in Table 13 a full comparison to state-of-the-art methods, including methods with LiDAR inputs and with post-processing such test-time training, Best-of-N scoring and ensembling. We note that the Best-of-N practice aims to score all predicted trajectories with ground truth and select the best trajectory based on the scores from the evaluation, which is not aligned with the evaluation protocol defined in NAVSIM [12, 5].
@@ Appendix H NAVSIM-v2 results before benchmark fix
From the table, we see that DrivoR ranks among the best-performing methods, with a much lighter model design (see more detail in Appendix E) and without extra training data. Also, DrivoR ranks first among state-of-the-art methods after the bug fix. This indicates that the errors made by DrivoR are mostly due to human errors, which are not penalized after the fix according to the NAVSIM-v2 benchmark [5].
@@
PDM-Closed [11] 94.4 98.8 100 99.5 100 93.5 99.3 97.7 36.0 88.1 90.6 96.3 98.5 100 83.1 73.7 91.5 25.4 51.3
Const. Vel. [12] 88.8 42.8 70.6 99.3 77.5 87.3 78.6 97.1 60.4 83.2 59.1 76.5 98.0 71.3 81.1 47.9 97.1 61.9 10.9
Ego Hist. MLP [12] 93.2 55.7 86.6 99.3 81.2 92.2 83.5 97.5 77.7 77.2 51.9 74.4 98.2 77.1 75.0 40.8 97.8 79.8 12.7
LTF [8] 96.2 79.6 99.1 99.6 84.1 95.1 94.2 97.6 79.1 77.8 70.2 84.3 98.1 85.1 75.7 45.4 95.8 76.0 23.1
RAP-DINO (ViT-H) [15] 97.1 94.4 98.8 99.8 83.9 96.9 94.7 96.4 66.2 83.2 83.9 87.4 98.0 86.9 80.4 52.3 95.2 52.4 36.9
GTRS-D (V2-99) [35] 98.7 91.4 95.8 89.2 99.4 94.4 99.3 98.8 72.8 69.5 98.7 90.1 95.1 54.6 96.9 94.1 40.4 49.7 41.7
GTRS-A (V2-99) [35] 98.9 87.9 95.1 88.8 99.2 89.6 99.6 98.8 76.1 80.3 99.1 86.0 94.7 53.5 97.6 97.1 54.2 56.1 42.1
@@
GTRS-D EVA-ViT-L [35] 97.6 95.8 99.8 99.0 77.2 97.8 95.3 97.3 46.7 91.9 91.3 92.7 99.0 72.7 90.4 53.8 94.1 41.6 43.4
GTRS-A (ViT-L) [35] 98.7 98.0 99.1 99.8 75.9 98.7 94.7 97.6 49.8 89.5 89.6 92.9 98.5 78.9 86.4 55.3 96.5 52.7 44.7
DriveSuprim (EVA-ViT-L) [53] 98.7 98.0 99.1 99.8 75.9 98.7 94.7 97.6 49.8 89.5 89.6 92.9 98.5 78.9 86.4 55.3 96.5 52.7 44.7
GTRS-D (ViT-L) [35] 98.9 98.2 99.8 99.6 73.9 98.9 95.3 97.3 40.0 91.5 90.8 94.7 98.5 70.8 90.1 55.4 97.2 54.2 45.3
DrivoR (ViT-S) 98.8 95.1 98.9 100 72.6 98.7 94.0 97.6 73.3 90.2 88.4 91.9 98.6 69.8 88.0 50.1 98.5 76.2 45.3
ZTRS (V2-99) [34] 98.9 97.6 100.0 100.0 66.7 98.9 96.2 96.7 44.0 91.1 90.4 95.8 99.0 63.6 89.8 60.4 97.6 66.1 45.5
@@
RAP-DINO (ViT-H) [15] 97.1 94.4 98.8 99.8 83.9 96.9 94.7 96.4 66.2 83.2 83.9 87.4 98.0 86.9 80.4 52.3 95.2 52.4 39.6
GTRS-D (V2-99) [35] 98.9 96.2 99.4 99.3 72.9 98.9 95.1 96.9 39.1 91.2 89.4 94.4 98.8 69.5 90.0 54.3 94.0 48.7 45.0
GTRS-A (V2-99) [35] 98.9 95.1 99.1 99.6 76.2 99.1 94.9 97.6 54.2 88.1 88.8 89.3 98.9 98.9 85.9 53.7 96.8 56.9 45.4
@@
ZTRS (V2-99) [34] 98.9 97.6 100 100 66.7 98.9 96.2 96.7 44.0 91.1 90.4 95.8 99.0 63.6 89.8 60.4 97.6 66.1 48.1
DrivoR (+65k SimScale data, ViT-S) 98.9 97.3 99.2 99.6 77.7 99.1 95.3 97.6 68.4 92.3 92.2 97.0 99.0 72.1 90.3 56.3 97.1 38.8 52.3
SimScale (+185k SimScale data, V2-99)  [45] 99.6 99.16 99.9 100 69.6 99.6 95.8 95.6 28.4 94.5 94.2 95.8 99.2 75.8 92.8 60.1 96.1 43.2 53.2
DrivoR (+134k SimScale data, ViT-S) 99.1 98.2 99.3 99.8 75.4 98.7 94.9 97.6 70.2 92.3 91.6 97.3 99.1 75.7 90.6 56.1 98.4 44.7 54.6